Given a single deck of cards, what is the probability of drawing either a king or a diamond? Hint – There are four kings in a deck and 13 diamonds. There are a total of 52 cards in a single deck.
Answer by checkley71(8403)

(4+12)/52=16/52=4/13 OR.3077 IS THE PROBABILITY OF DRAWING A KING OR A DIAMOND.
THERE ARE 4 KINGS (1 OF WHICH IS A DIAMOND)